# Corrupted but Not Broken: Rethinking the Impact of Corrupted Data in Visual Instruction Tuning

### v1

@@ 4.2 Analyzing the Superficial Impact of Corrupted Data
5 Defending against Corrupted Data
In Sec. 5.1, we show that post-training with a small set of clean data can effectively restore the performance of corrupted models. However, this raises an important question: How can we identify clean samples in a corrupted dataset? To answer this, in Sec. 5.2, we assess the ability of MLLMs to identify the cleanness of training samples and introduce self-validation as a robust solution and provide empirical understanding for it in Sec. 5.3. Finally, we combine post-training and self-validation as an effective corruption-robust training strategy in Sec. 5.4.
Figure 3: Average performance of the corrupted LLaVA-1.5 (LLaMA-3.1-8B) post-trained with an increasing number of clean data. 0 in x-axis indicates the original corrupted model without post-training.
@@ 5 Defending against Corrupted Data
Figure 4: Precision-recall curves of MLLM’s predictions on the correctness of 100K samples ($cr=50\%$). $x$-axis: recall; $y$-axis: precision. Solid and dotted line denote Val_PPL and PPL, respectively. Color represents the corruption ratio of the training dataset: $\bullet$0%, $\bullet$10%, $\bullet$20$\bullet$30%, $\bullet$40%, $\bullet$50%.
5.1 Post-training with Clean Data
In this section, we exploit the superficial impact of corrupted samples, as identified in Sec. 4.1, to restore the performance of corrupted models through post-training with clean data. The disabling of weights as discussed in Sec. 4.2 is another promising direction for performance recovery and we will leave it for future research. Specifically, for models fine-tuned on corrupted datasets with varying corruption ratios, we continue to fine-tune them with the same hyper-parameters as before. Note that here we use the original clean samples uniformly sampled from all datasets for post-training. However, we will introduce in Sec. 5.2 on how to obtain clean samples using the corrupted models.
Figure 3 shows the average evaluation performance for models fine-tune with varying corruption and post-trained with clean data at different sizes. As can be seen, even with only 20K clean samples, all models exhibit performance improvement after post-training. With more post-training data, the performance quickly converges to that of the clean model. This suggests post-training is a fast and cost-effective approach for performance recovery on corrupted models, without the need of re-training with clean data from scratch.
5.2 Can MLLMs Identify Corrupted Samples?
In this section, we show that MLLMs fine-tuned on corrupted training datasets can be used to identify corrupted samples. We use MLLMs that have been fine-tuned on datasets with corruption ratios from 0% to 50%, and run them on a dataset with 100K samples and $cr=50\%$. Each MLLM predicts a score (namely, Perplexity and Validation Perplexity to be presented in the following) on whether the response in each sample is corrupted or not. Samples with scores smaller than a threshold are considered clean. Details are in Appendix E.1.
The scores used are:
  1. 1.
    Perplexity (PPL): This quantifies how “surprised” or “uncertain” a model is when generating a response conditioned on an instruction (Marion et al., 2023). Samples with higher perplexities are considered as more likely to be corrupted. Note that PPL is proportional to the MLLM’s training loss (with their formulations in Appendix E.2) and treating large-loss samples as noisy samples in widely adopted in LNL (Han et al., 2018).
  2. 2.
    Validation Perplexity (Val_PPL): Using the following template, we directly prompt the corrupted model to predict whether a sample is corrupted (self-validation). The perplexity of the predicted word “No” is used as the score.
@@ 5.2 Can MLLMs Identify Corrupted Samples?
Effectiveness of PPL and Val_PPL.
Figure 4 compares the recalls and precisions of the scores across varying corruption levels. As can be seen, Val_PPL (solid curves) is effective in identifying clean data when trained with corrupted data. Specifically, at all corruption levels, these models (except the ones trained on clean data (black curves)) can achieve a precision of over 0.9 at a recall of 0.25. The reason why we focus on such a small recall is that Sec. 5.1 shows post-training with only a small amount of clean samples (e.g., 20K) can largely restore the model’s performance. This demonstrates the effectiveness of self-validation, which can be used in the proposed post-training approach. In contrast, PPL (dotted curves) performs worse. For example, at a corruption rate of 50% and recall of 0.25, PPL can only achieve a precision of less than 0.7. For the small-sized LLM (i.e., Qwen-2.5-0.5B), Val_PPL does not demonstrate much advantage. This suggests that self-validation is an emergent ability that can only be observed in larger language models.
Figure 5: Left: Distribution of classification probability $\hat{p}(z=1|x_{c},\tilde{x}_{y};\bm{\theta})$. Color represents corruption ratios of datasets the model is fine-tuned on: $\bullet$0%$\bullet$10%, $\bullet$20%. Middle: Mean of $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ and $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ of corrupted samples. Right: Mean ratios of $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ to $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ for clean and corrupted samples. All results here are obtained from LLaMA-3.1-8B on a dataset with 100K samples and $cr=50\%$.
How Robust are Scores under Corruption?
As shown in Figure 4, the Val_PPL curves (non-black solid curves) at different corruption ratios are consistent and close to each other. In contrast, PPL deteriorates sharply as the corruption ratio increases, which is expected since training with corrupted data explicitly lowers the perplexity of the MLLM on corrupted samples, making clean and corrupted data less distinguishable.
5.3 Analysis on Val_PPL
An intriguing observation from Figure 4 is that the solid black curves are always below the solid non-black curves, in other words, the corrupted model is even better than the clean model in identifying corrupted samples. We provide empirical analysis for this phenomenon in this section.
@@ 5.3 Analysis on Val_PPL
Recall that the Val_PPL template asks the MLLM to predict correctness of a sample. Essentially, the MLLM estimates $p(z=1|x_{c},\tilde{x}_{y})$, the ground-truth probability that response $\tilde{x}_{y}$ is correct, with its output probability $\hat{p}(z=1|x_{c},\tilde{x}_{y};\bm{\theta})$. Figure 5 (left) shows the distributions of $\hat{p}$ for models fine-tuned from datasets with different corruption levels. As can be seen, for predictions on clean samples (solid curves), corruption in the fine-tuning data has little effect on the output probability distribution. In contrast, significant distribution shift is observed on that of the corrupted samples (dashed curves).
To understand, we rewrite $p(z=1|x_{c},\tilde{x}_{y})$ as:
$p(z=1|x_{c},\tilde{x}_{y})-1\propto-\frac{p(\tilde{x}_{y}|x_{c},z=0)}{p(\tilde{x}_{y}|x_{c})}.$ (1)
Proof is in Appendix E.3. This shows that $p(z=1|x_{c},\tilde{x}_{y})-1$ is proportional to the negative of the ratio between $p(\tilde{x}_{y}|x_{c},z=0)$ and $p(\tilde{x}_{y}|x_{c})$. Similar to Val_PPL, $p(\tilde{x}_{y}|x_{c},z=0)$ can be obtained by explicitly prompting the MLLM as follows:
[…]
Figure 5 (middle) shows the mean of $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ and $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ at varying corruption levels for corrupted samples. We observe a sharp rise in $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ (blue) when transitioning from a clean model to one trained with slight corruption ($cr=10\%$), followed by a slower growth as the corruption increases further. In contrast, $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ (orange) exhibits a modest increase and remains consistently lower than the other probability. This discrepancy [^2] A symmetrical but less pronounced trend is observed for clean samples in Appendix E.4. results in a sharp increase in their ratio when the corruption is first introduced, followed by a gradual decline with additional corruption, which is confirmed in Figure 5 (right) that visualizes the mean of the ratios.
Impact of Ratio on Sample Separability.
The gap between the mean ratios for clean and corrupted samples reflects the separability of MLLMs on those samples using Val_PPL. As can be seen, this gap initially widens with small corruption, suggesting improved separability, but then narrows as higher corruption increases the likelihood of corrupted data, shrinking the ratio. This aligns with the trends in Figure 4, where slight corruption enhances classification while excessive corruption degrades it.
Asymmetric Growth of Two Probabilities.
The probability $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ for corrupted data increases rapidly with slight corruption, while $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ requires more corruptions to grow. This asymmetry stems from two factors. (i) Fine-tuning on corrupted data refines MLLM’s understanding of incorrect responses thereby sharply boosts $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$. (ii) $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ grows more slowly as fine-tuning optimizes both the clean and corrupted data, leading to competing objectives. In contrast, $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ is conditioned on $z=0$, evolves independently and increases more freely.
@@
Sample selection (post-training)
@@
Table 2: Comparisons of different corruption-robust strategies on LLaVA-1.5 (LLaMA-3.1-8B), where Avg. refers to the average performance on 11 benchmarks (normalized to 0-100). Best results are in bold, and the second best ones are underlined. Results for Qwen-2.5 series can be found in Tables 6, 7, and 8 in the appendix.
Figure 6: Average task performance of models post-trained (or fine-tuned from scratch) on data from different sources.
5.4 Post-training with Self-Validated Samples
The findings post-training in Sec. 5.1 and self-validation in Sec. 5.2 shed lights on an effective strategy when fine-tuned MLLM with corrupted datasets. In detail, we can first fine-tune an MLLM on corrupted datasets. Then we use self-validation to filter clean samples and finally conduct post-training on the fine-tuned MLLM with those filtered samples to obtain the final MLLM, which is used for evaluation.
6 Experiments
In this section, we evaluate the effectiveness of the corruption-robust training paradigm proposed in Sec. 5. Sec. 6.1 compares existing corruption-robust strategies to ours and Sec. 6.2 performs ablation studies of the post-training and self-validation components.
6.1 Comparisons with Existing Methods
We conduct experiments on the 11 benchmark datasets introduced before with $cr=50\%$, by using LLaVA-1.5 built on LLama-3.1-8B and Qwen-2.5 series models. We compare with the following baselines.
@@ 6.1 Comparisons with Existing Methods
    Noise-robust loss functions: By default, the MLLM is trained with the Cross-Entropy (CE) loss, denoted by None (CE). We also consider two noise-robust loss functions from Menon et al. (2019): the Generalized Cross-Entropy (GCE) loss (Zhang and Sabuncu, 2018), which combines the CE loss and mean absolute error (MAE) through the Box-Cox transformation (Box and Cox, 1964), as MAE has shown to have better robustness to label noise (Ghosh et al., 2017); and the Phuber Cross-Entropy loss (Menon et al., 2019) (Phuber CE), which incorporates gradient clipping into CE. Their formulations are provided in Appendix F.
@@
    Online sample selection methods: They focus on selecting clean samples during training. MentorNet (Jiang et al., 2018) identifies small-loss samples as clean. Co-teaching (Han et al., 2018) trains two networks and exchanges small-loss samples between them to avoid error accumulation. JoCoR (Wei et al., 2020) enforces agreement between networks to prevent biased selection. Their implementations are detailed in Appendix F.
  3. 3.
    Sample selection methods based on post-training: Different from online sample selection methods, the proposed PPL and Val_PPL methods select samples after the model is trained. Moreover, we also use several scores: EL2N and GradNorm, which measure the $\ell_{2}$-norm of the output error vector and the gradient, respectively (Paul et al., 2021); and Entropy (Coleman et al., 2020), which reflects uncertainty in the output probabilities. Formal definitions of these scores are provided in Appendix F.
@@ Results.
Table 2 compares the various corruption-robust strategies. As can be seen, Val_PPL significantly restores the performance of a corrupted model, improving it from 49.13 to 60.2 on average, where the clean model achieves 61.65. Moreover, Val_PPL outperforms all the baselines on average and achieves the best results on 8 out of 11 evaluation tasks.
6.2 Ablation Study
To assess the effectiveness of self-validation, we conduct post-training on LLaVA-1.5 fine-tuned with 50% corruption ratio [^3] Results for other corruption levels are in Figure 18 in the appendix. using “clean data" from various sources. (i) GT: Clean data with responses directly from the dataset. This can be regarded as the “best" clean data possible); (ii) PPL (we choose it as it is widely adopted in LNL); and (iii) the proposed Val_PPL. In addition, to demonstrate the advantage of post-training, we also experiment with (iv) Val_PPL(Scratch), which fine-tunes the model from scratch using the samples selected by Val_PPL rather than post-training.
Figure 6 shows the average task performance of these models post-trained (or fine-tuned from scratch) using “clean data” from various sources with different sizes. As can be seen, for the 3B and larger models, Val_PPL performs as well as GT across different data sizes, rapidly approaching the performance of the clean model. However, for Qwen-2.5-0.5B, only GT can restore the model performance. This is consistent with our observation in Figure 4 that the self-validation is an emergent ability for larger LLMs. Further, we find that Val_PPL(Scratch) is consistently outperformed by its post-trained counterparts, demonstrating the sample-efficiency of post-training.

### v2

@@ 4.2 Negative Effects of Data Corruption is Reversible
(a) PPL
(b) PPL and Val_PPL
@@
Figure 5: Precision-recall curves of MLLM’s predictions on the correctness of 100K samples ($cr=50\%$). $x$-axis: recall; $y$-axis: precision. Solid and dotted line denote predictions based on Val_PPL and PPL, respectively. Color represents the corruption ratio of the training dataset: $\bullet$0%, $\bullet$10%, $\bullet$20$\bullet$30%, $\bullet$40%, $\bullet$50%.
To recover the performance of the MLLM fine-tuned on corrupted data, we remove the model parameters affected by corrupted data. To achieve this goal, we first try to identify weights in the MLLM that are particularly responsible for generating corrupted responses. Specifically, following Wei et al. (2024), we select weights with top-$q\%$ influence scores on a corrupted dataset but remove those that overlap with weights with top-$p\%$ influence on a clean dataset. This ensures that only weights contributing specifically to corrupted samples are considered. We use the SNIP score  (Lee et al., 2019) to compute the influence, and consider an MLLM fine-tuned on 100K VIT data samples with corruption ratio $cr=60\%$. With different choices of $(p,q)$, we remove weights that are only related to the corrupted data and report the performance after removing these weights. For comparison, we include results from models fine-tuned with 100K and 40K clean VIT data samples (denoted by Clean and Clean(40K), respectively). See more details on in Appendix D.
Table 1 shows the performance on the evaluation tasks after removing weights related to corrupted data. By removing fewer than 1.4% of the parameters, the corrupted model can restore its performance from 39.1 to 57.1, which is already close to that (i.e., 59.3) of the model fine-tuned with 40K clean data (all clean data under a $cr$ of 60%).
Implications.
Notably, removing weights relevant to corrupted data aims to “delete” effects of corrupted data rather than “override” with new knowledge. This confirms that the corrupted MLLM does retain capabilities on evaluated tasks.
5 Underlying Capabilities of MLLMs Fine-tuned with Corrupted Data
Motivated by the restrictive and reversible nature of the negative impact of corrupted data as analyzed in Sections 4.1 and 4.2, we conjecture that the MLLMs fine-tuned with corrupted data may possess the underlying capacity to identify clean training samples and introduce a simple method called self-validation in Section 5.1.
By further fine-tuning corrupted MLLMs with samples selected by the proposed self-validation process, we demonstrate that the corrupted MLLM can recover from the negative impacts of corruption without any external annotations in Section 5.2.
5.1 Corrupted MLLM Can Detect Clean Samples from Corrupted Dataset
To identify clean and noisy samples from a corrupted VIT dataset, a straight-forward idea following existing works on learning with label noise is to consider the sample loss Jiang et al. (2018); Han et al. (2018): larger sample loss may indicate the corresponding sample is likely to contain wrong responses. That can be implemented using the perplexity score (Marion et al., 2023), as we can prove that it is proportional to the MLLM’s training loss (details in Appendix E.1). We term this score as PPL in this paper.
Figure 5(a) compares the recalls and precisions (see details in Appendix E.2) of using PPL from MLLM fine-tuned with varying corruption levels to identify clean and noisy samples. As can be seen, MLLMs fine-tuned with clean data (the black dotted line) can accurately identify clean samples by only keeping small-loss samples. For example, Qwen-2.5-3B achieves over 0.85 precision at a recall of 0.5. On the contrary, MLLMs fine-tuned with corrupted data (non-black dotted lines) achieve worse performance with increasing corruption levels. Specifically, all their precisions drop below 0.65 at a recall of 0.5. In other words, existing loss-based sample selection approaches are not robust to data corruption and cannot be simply extended to VIT.
5.1.1 Utilize Self-Validation to Activate Corrupted MLLM’s Capability
Given that the negative effect of corrupted data is restricted to specific response formats (Section 4.1) and that the corrupted MLLMs still retain capabilities on the evaluated tasks (Section 4.2), we propose to instruct the corrupted MLLM with different response format that is unseen during training. Specifically, using the following template, we directly prompt the MLLM to predict whether a sample is corrupted:
@@ 5.1.1 Utilize Self-Validation to Activate Corrupted MLLM’s Capability
The perplexity of the predicted word “No” is then used as the score. To distinguish from the perplexity score of responses, we refer the perplexity of word “No” in self-validation as Val_PPL. Contrary to PPL, samples with smaller Val_PPL score are considered more likely to be corrupted.
Self-Validation is robust against corrupted data.
Figure 5(b) compares the recalls and precisions of using the Val_PPL scores (solid curves) from MLLM fine-tuned with varying corruption levels to identify clean and corrupted samples. For reference, we also include the corresponding precision-recall curve of clean MLLMs previously shown in Figure 5(a) (black dotted curve). Remarkably, for MLLMs with 3B parameters and above, models fine-tuned with corrupted data (non-black solid lines) obtain similar precision-recall curves compared to that of the clean model using PPL. Further, even at increasing corruption levels, those curves are consistent and close to each other (except for Qwen-2.5-7B at $cr$ of 50%). This demonstrates that corrupted MLLMs can effectively distinguish clean and corrupted samples using Val_PPL.
However, we note that the advantage of Val_PPL is less significant for the small-sized LLM (Qwen-2.5-0.5B). This suggests that self-validation is an emergent ability that is only observed in larger language models.
Figure 6: Distribution of classification probability $\hat{p}(z=1|x_{c},\tilde{x}_{y};\bm{\theta})$. Color represents corruption ratios of datasets the model is trained on: $\bullet$0%$\bullet$10%, $\bullet$20%.
5.1.2 Why do Corrupted Data Improve the Performance of Self-Validation?
An intriguing observation from Figure 5(b) is that with self-validation, the corrupted model can identify corrupted samples even better than the clean model, as the solid black curves are always below the solid non-black curves. In this section, we provide an empirical analysis for this phenomenon.
@@ 5.1.2 Why do Corrupted Data Improve the Performance of Self-Validation?
Recall that the Val_PPL template asks the MLLM to predict correctness of a sample. Essentially, the MLLM estimates $p(z=1|x_{c},\tilde{x}_{y})$, the (ground-truth) probability that response $\tilde{x}_{y}$ is correct, with its output probability $\hat{p}(z=1|x_{c},\tilde{x}_{y};\bm{\theta})$. Figure 6 shows the distributions of $\hat{p}$ on the clean samples and corrupted samples from a corrupted VIT dataset that is also used to finetune the MLLM. We can see that corruption in the fine-tuning data has little effect on the output probability distribution for clean samples (solid curves). This results in robustness of Val_PPL at various corruption ratios and can be explained by the restrictive nature of corrupted data on MLLMs as the corrupted dataset does not contain self-validation instructions. In contrast, significant distribution shift is observed on that of the corrupted samples (dashed curves), which is the cause of improved performance of Val_PPL when fine-tuned with corrupted data. However, this cannot be explained by the conclusions drawn so far.
@@
Further fine-tuning
@@
Table 2: Comparisons of different corruption-robust strategies at a corruption ratio of 50% on LLaVA-1.5 (LLaMA-3.1-8B), where Avg. refers to the average results on 11 benchmarks (normalized to 0-100). Best results are in bold, and the second best are underlined. Results for Qwen-2.5 series are in Tables 6, 7, and 8 in the appendix.
Improved understanding of corrupted data
leads to distribution shift of $\hat{p}(z=1|x_{c},\tilde{x}_{y};\bm{\theta})$ on the corrupted samples (see Appendix E.3 for a detailed analysis). Although there are no explicit annotations indicating which samples are corrupted, the MLLM is still, to some extent, aware of the corrupted data during fine-tuning. This awareness stems from its prior pre-training on image-text pairs, which enables the model to learn correct image-text alignment. However, the PPL score cannot benefit from the improved understanding of corrupted data because it shares the same response format as the corrupted data and is thereby affected.
Figure 7: Average task performance of models further fine-tuned (or fine-tuned from scratch) on data from different sources. The corruption ratios are 20% (low) and 50% (high), results for other corruption levels are in Figure 19 in the appendix.
5.2 Corrupted MLLM can Recover from Data Corruption by Self-Validation
To further assess the quality of the samples selected from self-validation with the Val_PPL score, we investigate whether these samples can be used to recover the benchmark performance of MLLM previously fine-tuned on corrupted data. We conduct experiments on the 11 benchmark datasets introduced in Section 3 with the corruption ratio $cr=50\%$, by using LLaVA-1.5 built on LLama-3.1-8B and Qwen-2.5 series models. Specifically, we further fine-tune the MLLMs previously trained on corrupted data with the samples selected by from self-validation. By default, we choose samples with the highest 30% Val_PPL scores for further fine-tuning (i.e., 30K). In addition to using samples selected from self-validation with the Val_PPL score, we also consider using the following scores to select samples: (i) the PPL score proposed in Section 5.1; (ii) EL2N and GradNorm, which measure the $\ell_{2}$-norm of the output error vector and the gradient, respectively (Paul et al., 2021); and (iii) Entropy (Coleman et al., 2020), which reflects uncertainty in the output probabilities. See detailed formulae of these scores in Appendix F.3.
To better benchmark the performance of using self-validation for performance recovery, we also introduce the following baseline methods from traditional LNL (learning with noisy labels) methods:
@@ 5.2 Corrupted MLLM can Recover from Data Corruption by Self-Validation
    Noise-robust loss functions. By default, the MLLM is trained with the Cross-Entropy (CE) loss, denoted by None (CE). We consider two noise-robust loss functions from Menon et al. (2019): (i) Generalized Cross-Entropy (GCE) loss (Zhang and Sabuncu, 2018), that combines CE loss and mean absolute error (MAE), which is more robust to label noise (Ghosh et al., 2017), and (ii) the Phuber Cross-Entropy loss (Menon et al., 2019) (Phuber CE), which incorporates gradient clipping into CE. Their detailed formulae are in Appendix F.1.
@@
    Online sample selection methods, which focus on selecting clean samples during training. MentorNet (Jiang et al., 2018) identifies small-loss samples as clean. Co-teaching (Han et al., 2018) trains two networks and exchanges small-loss samples between them to avoid error accumulation. JoCoR (Wei et al., 2020) enforces agreement between networks to prevent biased selection. Their details are in Appendix F.2.
@@ Results.
Table 2 compares the performance of various corruption-robust strategies. By using the Val_PPL score to select clean VIT data samples for further fine-tuning, we can significantly restore the performance of a corrupted model, improving it from 49.13 to 60.2 on average, where the clean model achieves 61.65. We also outperform all existing baseline methods on average and achieves the best results on 8 out of 11 evaluation tasks, which further justify the high quality of VIT data samples selected by the Val_PPL score.
5.2.1 Analysis
To further assess the effectiveness of self-validation, we conduct further fine-tuning on LLaVA-1.5 fine-tuned with 20% and 50% corruption ratios, which cover low and high corruption in practical scenarios, [^2] Results for other corruption levels are in Figure 19 of the appendix. using “clean data" from various sources. (i) GT: Clean data with responses directly from the clean dataset. This can be regarded as the “best" clean data possible); (ii) PPL (we choose it rather than GradNorm or Entropy because it is widely adopted in LNL and we studied it in Section 5.1); and (iii) the proposed Val_PPL. In addition, we also experiment with (iv) Val_PPL(Scratch), which fine-tunes the model from scratch using the samples selected by Val_PPL rather than after it is fine-tuned on corrupted data.
Figure 7 shows the average task performance of these models further fine-tuned (or fine-tuned from scratch) using “clean data” from various sources with different sizes. As can be seen, with GT, most MLLMs could largely restore the performance using 20K samples, which indicates that given clean data, further fine-tuning can be a fast and cost-effective approach for performance recovery on corrupted MLLMs. Notably, both at low (20%) and high (50%) corruption ratios, Val_PPL performs as well as GT for the 3B and larger MLLMs, rapidly approaching the performance of the clean model. However, for Qwen-2.5-0.5B, only GT can restore the model performance. This is consistent with our observation in Figure 5(b) that the self-validation is an emergent ability for larger LLMs. On the contrary, though PPL could achieve comparable results with to Val_PPL at low corruption, it is less effective at high corruption, demonstrating the robustness of self-validation at various corruption ratios. Further, we find that Val_PPL(Scratch) is consistently outperformed by its further fine-tuning counterparts, demonstrating the sample-efficiency of further finetuning.